# Optimal Deterministic Rendezvous in Labeled Lines

### v1

@@ 4.3 Second Argument: In-Sync Rendezvous
[…]
@@ 4.3.1 Mismatched Ruling Sets Rendezvous
We first consider the case where one of the agents has computed part of a ruling set of higher distance ${R^{*}}$ than the other agent. As $R_{\alpha,L}$ and $R_{\beta,L}$ are powers of $4$, $R_{\alpha,L}>R_{\beta,L}$ implies that $R_{\alpha,L}\geq 4R_{\beta,L}$. This large difference between the two distances allows us to prove rendezvous by similar arguments as in the previous section (Lemma 4.3).
@@
Assume $\tau\leq 10D$ as otherwise Lemma 4.3 already yields the claim.
Let agent $\alpha$ have computed the ruling set of higher distance, i.e., $R_{\alpha,L}>R_{\beta,L}$ and $R_{\alpha,L}\geq D$. Note that $R_{\alpha,L}\leq L/16$, so $L\geq 16D$. In the first $L$ rounds of SearchingWalk(rounds $T_{\alpha,L}+4L$ to $T_{\alpha,L}+5L$), $\alpha$ moves to the ruling set node $r_{\alpha,L}$ at distance at most $12R_{\alpha,L}-$ 1 from $v_{\alpha}$. It then performs a ZWalk of parameter $8R_{\alpha,L}$. As a result, all the nodes within distance $8R_{\alpha,L}-(2R_{\alpha,L}-1)-D>5R_{\alpha,L}$ from $v_{\beta}$ are explored by $\alpha$ between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$. As for agent $\beta$, we know that it completed its discovery phase of parameter $L$ in round $T_{\beta,L}=T_{\alpha,L}+\tau\leq T_{\alpha,L}+L$. Therefore, between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$, $\beta$ is also executing SearchingWalk. In this algorithm, it stays within distance $(2R_{\beta,L}-1)+8R_{\beta,L}\leq(5/2)R_{\alpha,L}$ from $v_{\beta}$. As $\beta$ between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$ stays within an interval of nodes that gets fully explored by $\alpha$ during those rounds, by Proposition 3.2, the agents have achieved rendezvous by round $T_{\alpha,L}+6L$. The case $R_{\beta,L}>R_{\alpha,L}$ is similar, only introducing an extra $\tau\leq 10D\leq L$ to the number of rounds needed to achieve rendezvous. ∎
@@ 4.3.2 Common Ruling Set Rendezvous
Suppose that $R_{\alpha,L}=R_{\beta,L}\geq D$ and $r_{\alpha,L}=r_{\beta,L}$. Then ${T_{\mathsf{rdv}}}\leq T_{\alpha,L}+6L\in O(L)$.
@@
As before we assume that $\tau\leq 10D$, since Lemma 4.3 otherwise already yields the claim.
The agents both take at most $82R_{\alpha,L}-1\leq L/$ 8 rounds from the start of their respective SearchingWalk to reach the node $r_{\alpha,L}=r_{\beta,L}$. Therefore, between rounds $T_{\alpha,L}+4L+L/8$ and $T_{\alpha,L}+5L$, $\alpha$ is at $r_{\alpha,L}=r_{\beta,L}$. $\beta$ reaches this node in round $T_{\beta,L}+4L+L/8\leq T_{\alpha,L}+\tau+4L+L/8<T_{\alpha,L}+6L$ at the latest, so it necessarily meets $\alpha$ at $r_{\alpha,L}=r_{\beta,L}$ during the beginning of SearchingWalk. ∎
@@
Again let $\tau\leq 10D$ (Lemma 4.3), and note that $\tau\leq 10R_{\alpha,L}\leq L$.
[…]
Let $i$ be the first bit where $c_{\alpha,L}$ and $c_{\beta,L}$ differ. Let us assume that $\alpha$ has a $1$ at index $i$ of its color while $\beta$ has a 00, the opposite case being similar. Between round $T_{\alpha,L}+5L+(2i+1)\cdot 32R_{\alpha,L}$ and $T_{\alpha,L}+5L+(2i+3)\cdot 32R_{\alpha,L}$, agent $\alpha$ is executing $\textnormal{{{ZWalk}}}(R_{\alpha,L})$ twice in succession. In particular, $\alpha$ scans the entire distance-$8R_{\alpha,L}$ neighborhood of $r_{\alpha,L}$ between round $T_{\alpha,L}+5L+(2i+2)\cdot 32R_{\alpha,L}$ and $T_{\alpha,L}+5L+(2i+3)\cdot 32R_{\alpha,L}$. Recall that $\tau\leq 10D\leq 10R_{\alpha,L}$ and $T_{\beta,L}=T_{\alpha,L}+\tau$. Agent $\beta$ is staying at $r_{\beta,L}$ for $64R_{\alpha,L}$ rounds starting from round $T_{\beta,L}+5L+(2i+1)\cdot 32R_{\alpha,L}\leq T_{\alpha,L}+5L+(2i+1)\cdot 32R_{\alpha,L}+10R_{\alpha,L}$. Therefore, $\alpha$ necessarily finds $\beta$ during its second ZWalk for bit $i$, by Proposition 3.2. See Fig. 3 for an illustration with $i=3$ as first index where their colors differ. The bound on ${T_{\mathsf{rdv}}}$ simply follows from the runtime of SearchingWalk. ∎
@@ 4.4 Full Analysis
We now have analyzed all the cases necessary to show Theorem 1.1.
@@
Recall that $T_{\alpha,L}\in O(L)$ (Proposition 4.1). If the delay $\tau$ between the startup times of the two agents is $10D$ or larger, by Lemma 4.3, the agents reach rendezvous in $T_{\alpha,O(D)}+O(D)\in O(D)$ rounds (${T_{\mathsf{rdv}}}\in O(D)$). We assume a smaller delay $\tau$ in what follows.
By Lemma 4.4, we know that $R_{\alpha,L}\geq D$ and $R_{\beta,L}\geq D$ for $L>\Theta(D\log^{*}\ell_{\min})$. For $L$ a large enough $\Theta(D\log^{*}\ell_{\min})$, if during the loop iteration of parameter $L$ the agents settle on different values $R_{\alpha,L}\neq R_{\beta,L}$ for their ruling sets’ distances, they achieve rendezvous before round $T_{\alpha,L}+O(L)=O(L)=O(D\log^{*}\ell_{\min})$ by Lemma 4.5. If they have the same value $R_{\alpha,L}=R_{\beta,L}$, then they also achieve rendezvous before round $T_{\alpha,L}+O(L)=O(L)=O(D\log^{*}\ell_{\min})$, whether they go to the same ruling set node (Lemma 4.6) or different ones (Lemma 4.7). ∎
@@ 5 Corollary for Finite Graphs
In this section, we give simple extensions of our result for the infinite line to analogous path-like graphs like finite paths and cycles. Adapting our algorithm to also handle these graphs is only the matter of introducing some simple new behavior when an agent detects that the graph is finite (e.g., when it finds a node of degree $1$ for a finite path). We prove Corollary 1.2.
@@
If $n\in O(D\log^{*}\ell_{\min})$, then after $\Theta(\min(n,D\log^{*}\ell_{\min}))$ rounds, the two agents have discovered the whole graph. In particular, they must both be awake. For the finite line, that already implies that the agents achieved rendezvous by Proposition 3.2. For the cycle, the agents are guaranteed to meet once they have both reached the node of minimum ID, which must occur after $\Theta(\min(n,D\log^{*}\ell_{\min}))$ rounds (both agents must be awake by the time one of the two agent detects the cycle).
If $n\geq c\cdot D\log^{*}\ell_{\min}$ for a sufficiently big constant $c$, then the part of the graph that is explored by the two agents in $(c/8)\cdot D\log^{*}\ell_{\min}$ is indistinguishable from a segment from an infinite line. In that case, the agents meet in $O(D\log^{*}\ell_{\min})$ rounds from our analysis for the infinite line. ∎
On finite lines and the infinite half-line, the time to rendezvous also decreases with the distance with the distance of the agents to the nearest endpoint (node of degree $1$) of the line.
@@ Acknowledgments
We thank Sebastian Brandt for interesting discussions.

### v2

@@ 4.4 Second Argument: In-Sync Rendezvous
Recall that $R_{\alpha,L}$ and $R_{\beta,L}$ are the values $R^{*}$} computed by the two agents in the loop iteration of parameter $L$. $R_{\alpha,L}$ is the maximal value $R\in\{4^{i}\mid i\in\mathbb{N}\}\cap[0,L/16]$ such that for some node $u$ within distance $R$ from $v_{\alpha}$, the output of $u$ in $\textnormal{{{ES-ColPathRulingSet}}}(R)$ can be inferred from nodes within distance $L$ from $v_{\alpha}$. $R_{\alpha,L}=\bot$ indicates that no such value $R$ exists for the loop iteration $L$. $R_{\beta,L}$ is the same variable for agent $\beta$.
The output of a node $u$ in $\textnormal{{{ES-ColPathRulingSet}}}(R)$ includes a node from a (R,R1)(R,R-1)-ruling set within distance R1R-1 from $u$ (possibly $u$ itself), by Lemma˜3.5. In a loop iteration where $R_{\alpha,L}\neq\bot$, this means that $\alpha$ has learned about a node from the $(R_{\alpha,L},R_{\alpha,L}-1)$-ruling set computed by algorithm $\textnormal{{{ES-ColPathRulingSet}}}(R_{\alpha,L})$ at distance at most $2R_{\alpha,L}-1$ from $v_{\alpha}$. We denote $r_{\alpha,L}$ this node, and $c_{\alpha,L}$ the color it got imparted in $\textnormal{{{ES-ColPathRulingSet}}}(R_{\alpha,L})$.
@@
Let $v_{\min}\in N^{D}(v_{\alpha})\cup N^{D}(v_{\beta})$ be the node s.t. $\ell(v_{\min})=\ell_{\min}$ and $R_{D}=4^{\lceil\log_{4}(2D)\rceil}$. Note that $2D\leq R_{D}<8D$, $\operatorname{dist}(v_{\min},v_{\alpha})\leq 2D\leq R_{D}$, and $\operatorname{dist}(v_{\min},v_{\beta})\leq 2D\leq R_{D}$.
Consider lines 5 and 6 in Algorithm˜5. $R_{\alpha,L}$ is the largest value that satisfies the following three properties: $R_{\alpha,L}$ is a power of $4$, $R_{\alpha,L}\leq L/16$, and agent $\alpha$ has discovered enough of the line in its discovery phase of parameter $L$ to compute the output of at least one node within distance $R_{\alpha,L}$ from $v_{\alpha}$ in an execution of $\textnormal{{{ES-ColPathRulingSet}}}(R_{\alpha,L})$. $R_{\beta,L}$ is defined similarly for agent $\beta$. We show that all these properties hold for $R_{D}=4^{\lceil\log_{4}(2D)\rceil}$ with the lower bound on $L$ given as hypothesis, which implies the lemma.
@@ 4.4.1 Mismatched Ruling Sets Rendezvous
We first consider the case where one of the agents has computed part of a ruling set of higher distance $R^{*}$} than the other agent. As $R_{\alpha,L}$ and $R_{\beta,L}$ are powers of $4$, $R_{\alpha,L}>R_{\beta,L}$ implies that $R_{\alpha,L}\geq 4R_{\beta,L}$. This large difference between the two distances allows us to prove rendezvous by similar arguments as in the previous section (Lemma˜4.3).
@@
Assume $\tau\leq 10D$ as otherwise Lemma˜4.3 already yields the claim.
Let agent $\alpha$ have computed the ruling set of higher distance, i.e., $R_{\alpha,L}>R_{\beta,L}$ and $R_{\alpha,L}\geq D$. Note that $R_{\alpha,L}\leq L/16$, so $L\geq 16D\geq 8\tau/5$. In the first $L$ rounds of SearchingWalk (rounds $T_{\alpha,L}+4L$ to $T_{\alpha,L}+5L$), $\alpha$ moves to the ruling set node $r_{\alpha,L}$ at distance at most $2R_{\alpha,L}-1$ from $v_{\alpha}$. It then performs a ZWalk of parameter $8R_{\alpha,L}$. As a result, all the nodes within distance $8R_{\alpha,L}-(2R_{\alpha,L}-1)-D>5R_{\alpha,L}$ from $v_{\beta}$ are explored by $\alpha$ between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$.
Consider agent $\beta$ during the same rounds. We know that $\beta$ started its discovery phase of parameter $L$ in round $T_{\beta,L}=T_{\alpha,L}+\tau\leq T_{\alpha,L}+L$. Therefore, between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$, $\beta$ has finished its discovery phase of parameter $L$ and is either staying idle at $v_{\beta}$ or executing SearchingWalk. Either way, $\beta$ remains within distance $(2R_{\beta,L}-1)+8R_{\beta,L}\leq(5/2)R_{\alpha,L}$ from $v_{\beta}$ in those rounds.
Therefore, between round $T_{\alpha,L}+5L$ and $T_{\alpha,L}+6L$, $\beta$ stays within an interval of nodes while $\alpha$ fully explores that interval of nodes. By Proposition˜3.2, the agents achieve rendezvous by round $T_{\alpha,L}+6L$. The case $R_{\beta,L}>R_{\alpha,L}$ is similar, only introducing an extra $\tau\leq 10D\leq L$ to the number of rounds needed to achieve rendezvous. ∎
@@ 4.4.2 Common Ruling Set Rendezvous
Suppose that $R_{\alpha,L}=R_{\beta,L}\geq D$ and $r_{\alpha,L}=r_{\beta,L}$. Then ${T_{\mathsf{rdv}}}\leq T_{\alpha,L}+5L\in O(L)$.
@@
As before we assume that $\tau\leq 10D$, since Lemma˜4.3 otherwise already yields the claim. As $D\leq R_{\alpha,L}$, and $R_{\alpha,L}\leq L/16$, this implies $\tau\leq 5L/8$.
The agents both take at most $2R_{\alpha,L}-1\leq L/8$ rounds from the start of their respective SearchingWalk to reach the node $r_{\alpha,L}=r_{\beta,L}$. Therefore, between rounds $T_{\alpha,L}+4L+L/8$ and $T_{\alpha,L}+5L$, $\alpha$ is at $r_{\alpha,L}=r_{\beta,L}$. $\beta$ reaches this node by round $T_{\beta,L}+4L+L/8\leq T_{\alpha,L}+\tau+4L+L/8\leq T_{\alpha,L}+4L+3L/4$ at the latest. Thus the agents necessarily meet at $r_{\alpha,L}=r_{\beta,L}$ during the beginning of SearchingWalk before round $T_{\alpha,L}+5L$. ∎
@@
Again let $\tau\leq 10D$ (Lemma˜4.3), and note that $\tau\leq 10R_{\alpha,L}\leq L$.
Consider $c_{\alpha,L}$ and $c_{\beta,L}$ the colors assigned to the nodes $r_{\alpha,L}$ and $r_{\beta,L}$ by the execution of $\textnormal{{{ES-ColPathRulingSet}}}(R_{\alpha,L})$. As we have that $\operatorname{dist}(v_{\alpha},r_{\alpha,L})\leq 2R_{\alpha,L}-1$, $\operatorname{dist}(v_{\beta},r_{\beta,L})\leq 2R_{\alpha,L}-1$, and $\operatorname{dist}(v_{\alpha},v_{\beta})\leq D\leq R_{\alpha,L}$, it holds that nodes $r_{\alpha,L}$ and $r_{\beta,L}$ are at distance $\operatorname{dist}(r_{\alpha,L},r_{\beta,L})\leq 5R_{\alpha,L}-2$. Algorithm $\textnormal{{{ES-ColPathRulingSet}}}(R)$ computes a colored (R,R1)(R,R-1)-ruling set such that nodes of the ruling set get distinct colors when at distance 9R19R-1 or less from one another (Lemma˜3.5). Therefore, $c_{\alpha,L}\neq c_{\beta,L}$.
Let $i$ be the first bit where $c_{\alpha,L}$ and $c_{\beta,L}$ differ. Let us assume that $\alpha$ has a $1$ at index $i$ of its color while $\beta$ has a 0, the opposite case being similar. Between round $T_{\alpha,L}+5L+(2i+1)\cdot 32R_{\alpha,L}$ and $T_{\alpha,L}+5L+(2i+3)\cdot 32R_{\alpha,L}$, agent $\alpha$ is executing $\textnormal{{{ZWalk}}}(8R_{\alpha,L})$ twice in succession. In particular, $\alpha$ scans the entire distance-$8R_{\alpha,L}$ neighborhood of $r_{\alpha,L}$ between round $T_{\alpha,L}+5L+(2i+2)\cdot 32R_{\alpha,L}$ and $T_{\alpha,L}+5L+(2i+3)\cdot 32R_{\alpha,L}$. Recall that $\tau\leq 10D\leq 10R_{\alpha,L}$ and $T_{\beta,L}=T_{\alpha,L}+\tau$. Agent $\beta$ is staying at $r_{\beta,L}$ for $64R_{\alpha,L}$ rounds starting from round $T_{\beta,L}+5L+(2i+1)\cdot 32R_{\alpha,L}\leq T_{\alpha,L}+5L+(2i+1)\cdot 32R_{\alpha,L}+10R_{\alpha,L}$. Therefore, $\alpha$ necessarily finds $\beta$ during its second ZWalk for bit $i$, by Proposition˜3.2. See Fig.˜3 for an illustration with $i=3$ as first index where their colors differ. The bound on $T_{\mathsf{rdv}}$} simply follows from the runtime of SearchingWalk. ∎
@@ 4.5 Full Analysis
We now have analyzed all the cases necessary to show Theorem˜1.1.
@@
Recall that $T_{\alpha,L}\in O(L)$ (Proposition˜4.1). If the delay $\tau$ between the startup times of the two agents is $10D$ or larger, by Lemma˜4.3, the agents reach rendezvous in $T_{\alpha,O(D)}+O(D)\in O(D)$ rounds (${T_{\mathsf{rdv}}}\in O(D)$). We assume a smaller delay $\tau$ in what follows.
By Lemma˜4.4, we know that $R_{\alpha,L}\geq D$ and $R_{\beta,L}\geq D$ for $L>\Theta(D\log^{*}\ell_{\min})$. For $L$ a large enough $\Theta(D\log^{*}\ell_{\min})$, if during the loop iteration of parameter $L$ the agents settle on different values $R_{\alpha,L}\neq R_{\beta,L}$ for their ruling sets’ distances, they achieve rendezvous before round $T_{\alpha,L}+O(L)=O(L)=O(D\log^{*}\ell_{\min})$ by Lemma˜4.5. If they have the same value $R_{\alpha,L}=R_{\beta,L}$, then they also achieve rendezvous before round $T_{\alpha,L}+O(L)=O(L)=O(D\log^{*}\ell_{\min})$, whether they go to the same ruling set node (Lemma˜4.6) or different ones (Lemma˜4.7). ∎
@@ 5 Corollary for Finite Graphs
In this section, we give simple extensions of our result for the infinite line to analogous path-like graphs like finite paths and cycles. Adapting our algorithm to also handle these graphs is only the matter of introducing some simple new behavior when an agent detects that the graph is finite (e.g., when it finds a node of degree $1$ for a finite path). We prove Corollary˜1.2.
@@
Suppose $n\in O(D\log^{*}\ell_{\min})$. After $\Theta(\min(n,D\log^{*}\ell_{\min}))$ rounds, if rendezvous has not yet occurred, the two agents have discovered the entire graph. In particular, they are both awake. For the finite line, this already implies that the agents achieved rendezvous by Proposition˜3.2. For the cycle, the agents are guaranteed to meet once they have both reached the node of minimum ID, which must occur after $\Theta(\min(n,D\log^{*}\ell_{\min}))$ rounds (both agents must be awake by the time one of the two agents detects the cycle).
If $n\geq c\cdot D\log^{*}\ell_{\min}$ for a sufficiently large constant $c$, then the part of the graph that is explored by the two agents in $(c/8)\cdot D\log^{*}\ell_{\min}$ is indistinguishable from a segment from an infinite line. In that case, the agents meet in $O(D\log^{*}\ell_{\min})$ rounds from our analysis for the infinite line. ∎
On finite lines and the infinite half-line, the time to rendezvous also decreases with the distance of the agents to the nearest endpoint (node of degree $1$) of the line.
@@ Acknowledgments
We thank Sebastian Brandt for interesting discussions and anonymous reviewers for their valuable feedback. This research was conducted while Alexandre Nolin was at the CISPA Helmholtz Center for Information Security.